# On Uncertainty Calibration for Equivariant Functions

### v2

@@ Abstract
Data-sparse settings such as robotic manipulation, molecular physics, and galaxy morphology classification are some of the hardest domains for deep learning. For these problems, equivariant networks can help improve modeling across undersampled parts of the input space, and uncertainty estimation can guard against overconfidence. However, until now, the relationships between equivariance and model confidence, and more generally equivariance and model calibration, has yet to be studied. Since traditional classification and regression error terms show up in the definitions of calibration error, it is natural to suspect that previous work can be used to help understand the relationship between equivariance and calibration error. In this work, we present a theory relating equivariance to uncertainty estimation. By proving lower and upper bounds on uncertainty calibration errors (ECE and ENCE) under various equivariance conditions, we elucidate the generalization limits of equivariant models and illustrate how symmetry mismatch can result in miscalibration in both classification and regression. We complement our theoretical framework with numerical experiments that clarify the relationship between equivariance and uncertainty using a variety of real and simulated datasets, and we comment on trends with symmetry mismatch, group size, and aleatoric and epistemic uncertainties. Code is available at https://github.com/EdwardBerman/EquivariantUQBounds
@@ 1 Introduction
Equivariant neural networks are a class of neural networks that encode group symmetries into the structure of the network architecture so that the symmetries do not need to be learned from data. Understanding both model calibration and confidence is particularly useful in the data-sparse settings where equivariant neural networks tend to thrive, such as pick-and-place robotics tasks (Kalashnikov et al., 2018; Wang et al., 2022b; a; Fu et al., 2023; Huang et al., 2023; 2024b; 2024a; Wang, 2025), galaxy morphology classification (Pandya et al., 2023; 2025a), and molecular physics (Zou et al., 2023; Ramakrishnan et al., 2014). While equivariance has proved effective in these domains, it does have some drawbacks, including diminishing benefits at scale as in the case of extremely large datasets or aggressive augmentation (Wang et al., 2023b; Klee et al., 2023; Gruver et al., 2023; Brehmer et al., 2024; Abramson et al., 2024), provable degradation on model performance in cases of symmetry mismatch (Wang et al., 2024), more complex architectures, and higher compute costs.
@@
The purpose of this work is to address both the lack of a theory relating equivariance to uncertainty estimation, and the scarcity of experiments exploring this relationship in practice. To accomplish this, we extend the error bounds given by Wang et al. (2024) to a broader class of calibration losses. In this way, we can quantify the effect of equivariance not just on accuracy, but also on calibration. In particular, we show that calibration error is related to typical classification and regression errors over the fibers of each confidence prediction. These errors have known bounds for equivariant functions, which we use to provide lower and upper bounds on various calibration error metrics. We explore these threads in the context of correct, incorrect, and extrinsic (Wang et al., 2024) equivariance (the equivariance taxonomy hereafter), where correct means the ground truth function has the same symmetry as the equivariant model, incorrect means the ground truth function disagrees with the symmetry in the model, and extrinsic means the model symmetry transforms in-distribution data to out-of-distribution data. We also coin a metric, the aleatoric bleed, that prescribes how well a model is able to predict ground truth uncertainties themselves and lower bound it for equivariant models. This study illustrates that the effect of equivariance on model calibration is dependant on where we are in the equivariance taxonomy, supporting the previous line of work in Wang et al. (2023a; 2024). We supplement our bounds with experiments on a wide variety of real and simulated datasets that indicate how models can suffer from increased calibration error in instances of symmetry mismatch.
@@
    For the classification task, we provide bounds on calibration error and we tighten the bound in the limiting case of an invariance (Section˜4).
@@
    For the regression task, we generalize expected normalized calibration error beyond scalar values for mean and variance predictions. We derive its upper bound on certain equivariant models (Section˜5) and derive its lower bound in the special case of scalar-valued mean and variance predictions.
@@
    Additionally, we coin another metric, the aleatoric bleed, in order to study miscalibration in terms of aleatoric and epistemic uncertainty (Section˜6).
@@ 2 Related Work
Our work is closest thematically to Petrache & Trivedi (2023) and Wang et al. (2024), which both establish bounds on function generalization under various assumptions of symmetry mismatch, but neither study calibration errors specifically. We also consider the equivaraince taxonomy established in Wang et al. (2024), but depict its ramifications on uncertainty quantification. Throughout this paper, we assume access to universal $G$-equivariant models, guaranteed by prior work (Maron et al., 2019; Yarotsky, 2022). Yet, universality alone does not ensure reliable uncertainty estimates, and therefore our focus is on deriving error bounds for calibration error. Our work also uses the strategy of decomposing the input and output spaces in order to prove some bounds, which we accomplish through taking the quotient by a group. This is a strategy similarly employed in Sannai et al. (2021); Lawrence (2022); Petrache & Trivedi (2023); Wang et al. (2024). Theory on the relevant group representations for functions that output probabilities is partly addressed in Bloem-Reddy et al. (2020); Dobriban & Yu (2025), but no bounds on expected calibration error are presented. Reasoning about distributions in terms of invariants also has a rich history in deriving uninformative priors for Bayesian analysis – see Jaynes (1968). We build on this by studying how equivariance can affect the reliability of Bayesian methods. Specifically, we look at the ability to separate different types of uncertainty using evidential regression, including on multivariate distributions, generalizing some of the work of Van der Linden et al. (2025) who study equivariant model selection when trained to predict a univariate distribution. We also build off of Gelberg et al. (2024) by studying the behavior of Bayesian models under various equivariance constraints other than the weight space permutation invariances they consider.
@@
A longstanding goal in the computational sciences is to separate model (epistemic) uncertainties from (aleatoric) uncertainties inherent to the data (Ulmer et al., 2021; Hüllermeier & Waegeman, 2021; Osband et al., 2023; Fuchsgruber et al., 2024). Previous work underscores how this separation can be difficult to perform in practice. Techniques such as those presented in Kendall & Gal (2017); Amini et al. (2020) often fail to distinguish these uncertainties due to effects such as loss shaping (Ovadia et al., 2019; Valdenegro-Toro & Mori, 2022; Osband et al., 2023; Wimmer et al., 2023; Nevin et al., 2024; Jürgens et al., 2024). Our work explores this problem in the context of symmetry. In particular, we explore how the epistemic uncertainty – the uncertainty often quantified by calibration errors – can be confused with aleatoric uncertainty due to symmetry mismatch.
@@
For learning tasks with inherent uncertainty, it is natural to design a neural network that approximates a probability distribution rather than a single point estimation. There are many ways of doing this, such as with Bayesian Neural Networks (Kononenko, 1989), Epistemic Neural Networks (Osband et al., 2023), Normalizing Flows (Kobyzev et al., 2020; Papamakarios et al., 2021), or using the softmax function (Goodfellow et al., 2016) to learn a categorical distribution. These approaches are often computationally expensive to train and sample in practice. Thus, we often employ parameterization techniques to constrain neural networks to output simplified probability distributions and train them using a negative log likelihood loss derived from said distribution. For example, mean variance estimators (MVE) are the simplest type of neural network that predicts a parameterized distribution. Instead of predicting a single output, MVEs predict a mean $\mu$ and a variance $\sigma^{2}$ (Nix & Weigend, 1994; Seitzer et al., 2022). There is also work extending MVEs to learn covariances for multivariate distributions (Tomczak et al., 2020) and linear combinations of Gaussians (Diakonikolas et al., 2020). Amini et al. (2020) extend MVEs by imposing a prior on $\mu$ and $\sigma^{2}$ and performing evidential regression, which in turn provides enough parameters to disentangle aleatoric and epistemic uncertainties. A unique feature of our work is that we use models that predict parameterized distributions in order to define calibration error in a regression setting and also to experimentally test the affect of equivariance on the ability to learn a reliable uncertainty estimates.
@@
In classification, probabilistic models (e.g. logistic regression or softmax classifiers) output a distribution over labels, and the predicted label is chosen as the one with maximum probability. However, even if a model gives a label $y$ the highest probability $p$, that does not mean the model will necessarily be correct with probability $p$. This mismatch is often quantified with the expected calibration error (ECE) (Guo et al., 2017) and is often estimated using binning procedures that approximate the continuous push-forward density of different confidence regions. Miscalibration can analogously be measured for regression tasks (Pernot, 2023; Levi et al., 2022b). The idea is to compare true labels $y$ with a predicted mean $\mu$ and variance $\sigma^{2}$ of a MVE. One should expect the squared errors $(y-\mu)^{2}$ to average out to the variance $\sigma^{2}$. This idea was made precise by Levi et al. (2022a), who proposed the expected normalized calibration error (ENCE) to quantify this exact discrepancy. A key limitation of their work is that it is formulated in terms of binning approximations rather than in terms of a continuous probability density. Another key limitation of their work is that they assume mean and variance are scalar values. We generalize these metrics for continuous densities and multivariate normal distributions. Beyond ECE and ENCE, some works propose calibration objectives and training procedures in terms of coverage (Gneiting & Raftery, 2007; Lemos et al., 2023; Sun et al., 2023), distributional calibration (Kuleshov et al., 2018), and post-hoc variance scaling (Laves et al., 2020). We focus on ECE and ENCE because they are clearly formulated objectives that can be studied independently of the training process itself. The work of Sun et al. (2023) suggests that equivariance can improve model calibration, but a theoretical justification for this is not present in the literature and is something we comment on in this work.
@@ 3 Background
We review the definition of equivariance and how symmetry constraints in models may be mismatched with the datasets they are applied to. Additionally, we review evidential regression, a technique for learning model and data uncertainties as distinct outputs of a neural network.
@@ 3.1 Equivariance
Here, we give precise definitions of equivariance and invariance. For a general review of the mathematical background, we direct the reader to Artin (1998); Esteves (2020); Hall (2013).
Let $\phi:\mathcal{X}\rightarrow\mathcal{Y}$ be a map between input and output vector spaces $\mathcal{X}$ and $\mathcal{Y}$. Let $G$ be group with representations $\rho^{\mathcal{X}}$ and $\rho^{\mathcal{Y}}$ which transform vectors in $\mathcal{X}$ and $\mathcal{Y}$ respectively. Representations map group elements to invertible linear transformations. When clear, we omit the representation map and write $gx$ for $\rho^{\mathcal{X}}(g)x$. The map $\phi:\mathcal{X}\rightarrow\mathcal{Y}$ is equivariant if
$\rho^{\mathcal{Y}}(g)[\phi(x)]=\phi(\rho^{\mathcal{X}}(g)[x])\ ,\ \text{for all }g\in G,x\in\mathcal{X}\ .$
Invariance is a special case of equivariance in which $\rho^{\mathcal{Y}}=\rm{Id}^{\mathcal{Y}}$ for all $g\in G$. I.e., a map $\phi:\mathcal{X}\rightarrow\mathcal{Y}$ is invariant if it satisfies
$\phi(x)=\phi(\rho^{\mathcal{X}}(g)[x])\ ,\ \text{for all }g\in G,x\in\mathcal{X}.$
@@ 3.2 Equivariant Learning
Aligning with Wang et al. (2024), we first establish an equivariant learning framework that describes assumptions on the ground truth and hypothesis class.
Equivariant Learning Framework.
We work in the deterministic realizable case of statistical learning theory: data distribution is defined on $X$ and given by the probability density function $p:X\to\mathbb{R}$, labels are given deterministically by a ground truth function $f:X\to Y$. Following this standard statistical learning framework, we assume that training and testing samples are drawn i.i.d. from the same underlying distribution $p$, i.e., no distribution shift occurs at test time. The goal for a function space $\mathscr{H}=\{h:X\to Y\}$ is to fit the function $f$ by minimizing an error function $\text{err}(h)$. Let $\mathds{1}(x)$ be an indicator function that equals to $1$ if the condition in the argument is satisfied and 0 otherwise. In classification, $\text{err}(h)$ is the classification error rate; for regression tasks, the error function is a $L_{2}$ norm function,
@@
Wang et al. (2024) establish a taxonomy which describes the relationship of the symmetry of the function space to the symmetry in the data. We review the definitions of correct, incorrect, and extrinsic equivariance from Wang et al. (2024) . These definitions help us understand the ability of equivariant functions to approximate datasets that may or may not have the same symmetries. A key inclusion here is extrinsic symmetry, which describes the case where the action of the group moves data points out of the support of their original distribution.
@@
For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)\neq 0$ and $f(gx)=gf(x)$, $h$ has correct equivariance with respect to $f$ at $x$ under transformation $g$. For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)\neq 0$ and $f(gx)\neq gf(x)$, $h$ has incorrect equivariance with respect to $f$ at $x$ under transformation $g$. For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)=0$, $h$ has extrinsic equivariance with respect to $f$ at $x$ under transformation $g$.
If $f$ has point-wise correct equivariance for all $x\in X$, $g\in G$, then we simplify and say that $f$ has correct equivariance. The nomenclature follows for incorrect and extrinsic equivariance as well. In the case that $f$ has correct equivariance, we assume $f$ lies in the considered hypothesis class $\mathscr{H}$. This assumption aligns with Wang et al. (2024) and is realistic, since several universality results for equivariant models are already known (Yarotsky, 2022).

### v4

@@ Abstract
Data-sparse settings such as robotic manipulation, molecular physics, and galaxy morphology classification are some of the hardest domains for deep learning. For these problems, equivariant networks can help improve modeling across undersampled parts of the input space, and uncertainty estimation can guard against overconfidence. However, until now, the relationships between equivariance and model confidence, and more generally equivariance and model calibration, have yet to be studied. Since traditional classification and regression error terms show up in the definitions of calibration error, it is natural to suspect that previous work can be used to help understand the relationship between equivariance and calibration error. In this work, we present a theory relating equivariance to uncertainty estimation. By proving lower and upper bounds on uncertainty calibration errors (ECE and ENCE) under various equivariance conditions, we elucidate the generalization limits of equivariant models and illustrate how symmetry mismatch can result in miscalibration in both classification and regression settings. We complement our theoretical framework with numerical experiments that clarify the relationship between equivariance and uncertainty using a variety of real and simulated datasets, and we comment on trends with symmetry mismatch, group size, and aleatoric and epistemic uncertainties.
@@ 1 Introduction
Equivariant neural networks are a class of neural networks that encode group symmetries into the structure of the network architecture so that the symmetries do not need to be learned from data. Understanding both model calibration and confidence is particularly useful in the data-sparse settings where equivariant neural networks tend to thrive, such as pick-and-place robotics tasks (Kalashnikov et al., 2018; Wang et al., 2022b; a; Fu et al., 2023; Huang et al., 2023; 2024b; 2024a; Wang,Dian, 2025), galaxy morphology classification (Pandya et al., 2023; 2025a), and molecular physics (Zou et al., 2023; Ramakrishnan et al., 2014). While equivariance has proved effective in these domains, it does have some drawbacks, including diminishing benefits at scale as in the case of extremely large datasets or aggressive augmentation (Wang et al., 2023b; Klee et al., 2023; Gruver et al., 2023; Brehmer et al., 2024; Abramson et al., 2024), provable degradation on model performance in cases of symmetry mismatch (Wang et al., 2024), more complex architectures, and higher compute costs.
@@
The purpose of this work is to address both the lack of a theory relating equivariance to uncertainty estimation, and the scarcity of experiments exploring this relationship in practice. To accomplish this, we extend the error bounds given by Wang et al. (2024) to a broader class of calibration losses. In this way, we can quantify the effect of equivariance not just on accuracy, but also on calibration. In particular, we show that calibration error is related to typical classification and regression errors over the preimage of each confidence prediction. These errors have known bounds for equivariant functions, which we use to provide lower and upper bounds on various calibration error metrics. We explore these threads in the context of the equivariance taxonomy from (Wang et al., 2024), which describes scenarios when the ground truth function has the same symmetry as the equivariant model, the ground truth function disagrees with the symmetry in the model, and the model symmetry transforms in-distribution data to out-of-distribution data. We also define a new metric, the aleatoric bleed, that prescribes how well a model is able to distinguish between different types of uncertainty. We provide a lower bound on aleatoric bleed for equivariant models. Our study illustrates that the effect of equivariance on model calibration is dependent on the extent of the symmetry mismatch, a finding similarly reported for accuracy in Wang et al. (2023a; 2024). We also perform experiments on a wide variety of real and simulated datasets, empirically showing that symmetry mismatch can increase calibration error and demonstrating that our theoretical findings are informative in practical settings.
@@
    For classification, we provide bounds on calibration error and we tighten the bound in the limiting case of an invariance (section˜4).
@@
    For regression, we generalize expected normalized calibration error beyond scalar values for mean and variance predictions. We derive its upper bound on certain equivariant models (section˜5) and derive its lower bound in the special case of scalar-valued mean and variance predictions.
@@
    Additionally, we define a new metric, the aleatoric bleed, which quantifies miscalibration in terms of aleatoric and epistemic uncertainty (section˜6).
@@ 2 Related Work
Our work studies the problem of symmetry mismatch between model and data through the lens of approximation error. We are closest thematically to Petrache and Trivedi (2023) and Wang et al. (2024), which both establish bounds on function generalization under various assumptions of symmetry mismatch, but neither study calibration errors specifically. We also consider the equivariance taxonomy established in Wang et al. (2024), but depict its ramifications on uncertainty quantification. Throughout this paper, we assume access to universal $G$-equivariant models, guaranteed by prior work (Maron et al., 2019; Yarotsky, 2022). Yet, universality alone does not ensure reliable uncertainty estimates, and therefore our focus is on deriving error bounds for calibration error. In order to prove these bounds, our work also uses the strategy of decomposing the input and output spaces, which we accomplish through taking the quotient by the symmetry group. This is a strategy similarly employed in Sannai et al. (2021), Lawrence (2022), Petrache and Trivedi (2023), and Wang et al. (2024). Theory on the relevant group representations for functions that output probabilities is partly addressed in Bloem-Reddy et al. (2020) and Dobriban and Yu (2025), but no bounds on expected calibration error are presented. Reasoning about distributions in terms of invariants also has a rich history in deriving uninformative priors for Bayesian analysis – see Jaynes (1968). We build on this by studying how equivariance can affect the reliability of Bayesian methods. Specifically, we look at the ability to separate different types of uncertainty using evidential regression, including on multivariate distributions, generalizing some of the work of Van der Linden et al. (2025) who study equivariant model selection when trained to predict a univariate distribution. We also build off of Gelberg et al. (2024) by studying the behavior of Bayesian models under various equivariance constraints other than the weight space permutation invariances they consider.
@@
A longstanding goal in the computational sciences is to separate model (epistemic) uncertainties from (aleatoric) uncertainties inherent to the data (Ulmer et al., 2021; Hüllermeier and Waegeman, 2021; Osband et al., 2023; Fuchsgruber et al., 2024; Chau et al., 2025). Previous work underscores how this separation can be difficult to perform in practice. Techniques such as those presented in Kendall and Gal (2017); Amini et al. (2020) often fail to distinguish these uncertainties due to effects such as loss shaping (Ovadia et al., 2019; Valdenegro-Toro and Mori, 2022; Osband et al., 2023; Wimmer et al., 2023; Nevin et al., 2024; Jürgens et al., 2024). Our work explores this problem in the context of symmetry. In particular, we explore how the epistemic uncertainty—the uncertainty often quantified by calibration errors—can be confused with aleatoric uncertainty due to symmetry mismatch.
@@
For learning tasks with inherent uncertainty, it is natural to design a neural network that approximates a probability distribution rather than a single point estimation. There are many ways of doing this, such as with Bayesian Neural Networks (Kononenko, 1989), Epistemic Neural Networks (Osband et al., 2023), Normalizing Flows (Rezende and Mohamed, 2015; Kobyzev et al., 2020; Papamakarios et al., 2021), or using the softmax function (Goodfellow et al., 2016) to learn a categorical distribution. These approaches are often computationally expensive to train and sample in practice. Thus, we often employ parameterization techniques to constrain neural networks to output simplified probability distributions and train them using a negative log likelihood loss derived from said distribution. For example, mean variance estimators (MVE) are the simplest type of neural network that predicts a parameterized distribution. Instead of predicting a single output, MVEs predict a mean $\mu$ and a variance $\sigma^{2}$ (Nix and Weigend, 1994; Seitzer et al., 2022). There is also work extending MVEs to learn covariances for multivariate distributions (Tomczak et al., 2020) and linear combinations of Gaussians (Diakonikolas et al., 2020). Amini et al. (2020) extend MVEs by imposing a prior on $\mu$ and $\sigma^{2}$ and performing evidential regression, which in-turn provides enough parameters to disentangle aleatoric and epistemic uncertainties. A unique feature of our work is that we use models that predict parameterized distributions in order to define calibration error in a regression setting and also to experimentally test the effect of equivariance on the ability to learn a reliable uncertainty estimate.
@@
In classification, probabilistic models (e.g. logistic regression or softmax classifiers) output a distribution over labels, and the predicted label is chosen as the one with maximum probability. However, even if a model gives a label $y$ the highest probability $p$, that does not mean the model will necessarily be correct with probability $p$. This mismatch is often quantified with the expected calibration error (ECE) (Guo et al., 2017) and is often estimated using binning procedures that approximate the continuous push-forward density of different confidence regions. Miscalibration can analogously be measured for regression tasks (Pernot, 2023; Levi et al., 2022b). The idea is to compare true labels $y$ with a predicted mean $\mu$ and variance $\sigma^{2}$ of a MVE. One should expect the squared errors $(y-\mu)^{2}$ to average out to the variance $\sigma^{2}$. This idea was made precise by Levi et al. (2022a), who proposed the expected normalized calibration error (ENCE) to quantify this exact discrepancy. A key limitation of their work is that it is formulated in terms of binning approximations rather than in terms of a continuous probability density. Another key limitation of their work is that they assume mean and variance are scalar values. We generalize these metrics for continuous densities and multivariate normal distributions. Beyond ECE and ENCE, some works propose calibration objectives and training procedures in terms of coverage (Gneiting and Raftery, 2007; Lemos et al., 2023; Sun et al., 2023), distributional calibration (Kuleshov et al., 2018), and post-hoc variance scaling (Laves et al., 2020). We focus on ECE and ENCE because they are clearly formulated objectives that can be studied independently of the training process itself. The work of Sun et al. (2023) suggests that equivariance can improve model calibration, but a theoretical justification for this is not present in the literature and is something we comment on in this work.
@@ 3 Background
We review the definition of equivariance and how symmetry constraints of a model class may be mismatched with a given dataset. Additionally, we review evidential regression, a technique for learning model and data uncertainties as distinct outputs of a neural network.
@@ 3.1 Equivariance
Here, we give precise definitions of equivariance and invariance. For a general review of the mathematical background, we direct the reader to Artin (1998); Hall (2013); Esteves (2020).
Let $\phi:X\rightarrow Y$ be a map between input and output vector spaces $X$ and $Y$. Let $G$ be a group with representations $\rho^{X}$ and $\rho^{Y}$ which transform vectors in $X$ and $Y$ respectively. Representations are group homomorphisms which map group elements to invertible linear transformations. When clear, we omit the representation map and write $gx$ for $\rho^{X}(g)x$. The map $\phi:X\rightarrow Y$ is equivariant if
$\rho^{Y}(g)\phi(x)=\phi(\rho^{X}(g)x)\ ,\ \text{for all }g\in G,x\in X\ .$
Invariance is a special case of equivariance in which $\rho^{Y}=\rm{Id}^{Y}$ for all $g\in G$. I.e., a map $\phi:X\rightarrow Y$ is invariant if it satisfies
$\phi(x)=\phi(\rho^{X}(g)[x])\ ,\ \text{for all }g\in G,x\in X.$
@@
In the following, we give a simple example of fundamental domain.
Example 1.
Let $G=\mathrm{SO}(n)$, $X=\mathbb{R}^{n}$ and $F=\left\{(x,0,\dots,0)\mid x\in\mathbb{R}_{\geq 0}\right\}$. The closed set $F$ is a fundamental domain of $\mathrm{SO}(n)$ in $\mathbb{R}^{n}$. Indeed, the intersection $g_{1}F\cap g_{2}F=\{0\}$ has measure 0 for each distinct $g_{1},g_{2}\in G$ and $\mathbb{R}^{n}=\cup_{g\in\mathrm{SO}(n)}gF$.
@@ 3.2 Equivariant Learning
Following Wang et al. (2024), we first establish an equivariant learning setting that describes assumptions on the ground truth and hypothesis class. We work in the deterministic realizable case of statistical learning theory: data distribution is defined on $X$ and given by the probability density function $p\colon X\to\mathbb{R}$, labels are given deterministically by a ground truth function $f\colon X\to Y$. Following this standard statistical learning framework, we assume that training and testing samples are drawn i.i.d. from the same underlying distribution $p$, i.e., no distribution shift occurs at test time. The goal for a function space $\mathscr{H}=\{h\colon X\to Y\}$ is to fit the function $f$ by minimizing an error function $\text{err}(h)$. Let $\mathds{1}(x)$ be an indicator function that equals $1$ if the condition in the argument is satisfied and 0 otherwise. In classification, $\text{err}(h)$ is the classification error rate; for regression tasks, the error function is a $L_{2}$ norm function,
@@
Wang et al. (2024) establish a taxonomy which describes the relationship of the symmetry of the function space to the symmetry in the data. We review their definitions of correct, incorrect, and extrinsic equivariance. These definitions help us understand the ability of equivariant functions to approximate datasets that may or may not have the same symmetries. An important part of this taxonomy is extrinsic symmetry, which describes the case where the action of the group moves data points out of the support of their original distribution.
@@
Assume $h$ is equivariant with respect to a group $G$. For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)\neq 0$ and $f(gx)=gf(x)$, $h$ has correct equivariance with respect to $f$ at $x$ under transformation $g$. For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)\neq 0$ and $f(gx)\neq gf(x)$, $h$ has incorrect equivariance with respect to $f$ at $x$ under transformation $g$. For $g\in G$ and $x\in X$ where $p(x)\neq 0$, if $p(gx)=0$, $h$ has extrinsic equivariance with respect to $f$ at $x$ under transformation $g$.
If $f$ has point-wise correct equivariance for all $x\in X$, $g\in G$, then we say $f$ has correct equivariance. The same follows for incorrect and extrinsic equivariance as well. In the case that $f$ has correct equivariance, we assume $f$ lies in the considered hypothesis class $\mathscr{H}$. This assumption aligns with Wang et al. (2024) and is realistic, since several universality results for equivariant models are already known (Yarotsky, 2022).